# The Value of Information in Human-AI Decision-making

### v3

@@ 3.3 Robustness Analysis of Information Order
Our approach assumes the specification of a decision problem on which agents’ decisions and use of information are evaluated. However, ambiguity around the appropriate decision problem, and in particular, the appropriate scoring rule, is not uncommon in human-AI decision settings. Ambiguity can arise as a result of challenges in eliciting utility functions and/or variance in these functions across decision-makers or groups of instances; for example, doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). Blackwell’s comparison of signals (Blackwell et al., 1951) is an appropriate tool for addressing ambiguity about the payoff function, as it defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks. (Li et al., 2022; Kleinberg et al., 2023).
@@
Extending this to agent-complementary information value, we say that $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
@@
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Section 4.
4 Experiment I: Model Comparison on Chest Radiograph Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025). We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
4.1 Data and Model
We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023). The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. (2019) and Heidenreich et al. (2022). We label the radiology reports by a rule-based tool (Irvin et al., 2019) and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports. The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$. We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017). Our training set contains 12,228 images, and the validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models is: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, and VisionTransformer with $0.80$.
We construct the scale of performances by a no-information bound and a full-information bound. The no-information bound is $\mathrm{R}^{\pi,S}(\emptyset)$, the baseline as we define the information value. The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models. We do not consider the radiology images as a basic signal in this task because of the problem of computational feasibility with high-dimensional signals. We use the five vision models’ predictions as embeddings of the original images. See more detail on the algorithm of choosing the optimal embeddings in Appendix B.
4.2 Results
Can the AI models complement human judgment?
5 Experiment II: Behavioral Analysis on Deepfake Detection
We apply our framework to analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
Figure 2: Information value calculated under our framework in the information model defined by the experiment of Groh et al. (2022). Basic signals include the seven video level features and three types of agent decisions. The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
5.1 Data and Model
We define the information model on the experiment data of Groh et al. (2022). Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model. Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision. This produced a final decision ($D^{\text{Human-AI}}$). Both decisions were elicited as a percentage indicating how confident the participant was that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI model to the same 100-scale probability scale available to study participants.
We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from the mean absolute error used by Groh et al. (2022), but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. (2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Section 4 while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
5.2 Results
How much decision-relevant information do agents’ decisions offer?
We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$). This is consistent with the findings of Guo et al. (2024) that humans are bad at distinguishing when AI predictions are correct.
6 Experiment III: Information-based Local Explanation on Recidivism Prediction
Figure 3: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation, where each point represents an instance from the test set. (C-F) Waterfall plots of applying SHAP and ILIV-SHAP on instance 4 and instance 18. The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output). See further comparisons between SHAP and ILIV-SHAP in Appendix D.
6.1 Data and Model
We use the COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history. We merge the dataset with the experimental data from Lin et al. (2020) to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al. (2022)’s experiment on laypeople. Human decisions were elicited on a 30-point probability scale. Merging the two datasets produces 9,269 instances (defendants). For the AI model, we trained an extreme gradient boosting (XGBoost) model (Chen and Guestrin, 2016) on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances. We round the model predictions to the same 30-point probability scale available to study participants in Lin et al. (2020).
We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism (Green and Chen, 2019a; Lin et al., 2020; Fogliato et al., 2021). The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years. We use the features of the defendant contained in the COMPAS dataset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrests and charges for several offense categories), and the predicted score from the COMPAS system [^4] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx. We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
6.2 Results
@@ 7 Discussion and Limitations
We propose a decision-theoretic framework for assigning value to information in human-AI decision-making. Our methods quantify the additional information value of any signals over an agent’s decisions, and can be applied to any combination of artificial or human agent judgments. The three demonstrations we provide show how quantified information value can drive model selection, empirical evaluation, and explanation design. These are just a few of many possible use cases. For example, explanation approaches could be compared in terms of their information value to human decisions. Information value analysis could also drive elicitation of human private signals upon identifying that human decisions contain AI-complementary information. New explanation strategies based on ILIV-SHAP could more directly integrate visualized information value with conventional depictions of feature importance.
It is important to note that when using ACIV to understand complementary information over human decisions, it cannot be interpreted as definitively establishing that particular signals were used by the human. It is always possible that the human has other private signals offering equivalent information to a feature being analyzed. This raises a question of whether it is definitive about use if it is an AI as the agent instead of a human. However, quantifying the value of observed information is an important first step toward learning about information that may exist beyond an explicit problem definition.
Our framework quantifies the best-attainable performance improvement from integrating signals in decisions. This does not necessarily mean that highlighting or otherwise emphasizing those signals for humans will lead to them performing better. However, the basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021). For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the rational decision-maker to only make monotonic decisions with AI prediction and their own predictions.
@@ Appendix A The Combinatorial Nature of the Value of Signals
We model the information value of a single signal over the existing information in agent decisions. When decision-makers are provided with multiple signals, they might use them in combination. Therefore, our definition of information value in Definition 3.2 may overlook the signals’ value in combination with other signals. Signals can be complemented [Chen and Waggoner, 2016], i.e, they contain no information value by themselves but a considerable value when combined with other signals. For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ ight be uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$. In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff. To consider this complementation between signals, we use the Shapley value $\phi$ [Shapley, 1953] to interpret the contribution to information gain of each basic signal. The Shapley value calculates the average of the marginal contribution of a basic signal $\mathbf{\Sigma}_{i}$ in every combination of signals.
$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(\operatorname{ACIV}(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (10)
The Shapley value suggests how much information value of the basic signal is unexploited by the human decision-maker on average in all combinations.
@@
Appendix B Embeddings for High-dimensional Signals in Decision Problem
@@
Suppose the data-generating process is estimated from a set of observations $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
$\hat{\pi}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
In this case, the sample size $T$ that is required to get a good estimate of DGP is the exponential of the number of dimensions of the signal $v$. The high-dimensional signals, such as images and text, that are usual in human-AI decision problems are not computationally feasible to be quantified under our framework. We propose a systematic procedure to pick the optimal reduction model to represent the high-dimensional signal space in a low-dimensional embedding space.
The optimal reduction model and its generated embeddings should fulfill two properties. First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals. Second, the derived DGP that is implied by the embedding signals should be feasible to be estimated from the observed dataset. We quantify the above objectives by a framework based on Train/Test split.
Algorithm 2 Embedding high-dimensional signals in decision problems
1:  Input: Observed dataset $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Candidate model generator $C$.
2:  $R^{*}=0$

### v4

@@ 3.3 Robust Analysis of Information Order
Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem. However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule. In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). Blackwell’s comparison of signals (Blackwell et al., 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks. (Li et al., 2022; Kleinberg et al., 2023).
@@
Extending this to ACIV, $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
@@
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix D.
4 Demonstrations
We demonstrate uses of our framework on three real-world decision-making tasks: medical diagnosis, deepfake video detection, and recidivism prediction.
First, we demonstrate the framework’s utility in model selection by evaluating how well different cardiac dysfunction prediction models trained on chest X-rays complement radiologists’ predictions (Rajpurkar et al., 2018; Johnson et al., 2019), showing how even among models with similar accuracy,some models strictly offer more complementary information than others over the radiologists. Next, we show how the framework can be used to evaluate how well humans exploit available features in a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022), how this compares to how an AI models exploits the information, and how providing AI assistance affects the human use. Suprisingly, even though AI assistance helps human-AI teams achieve higher decision accuracy, the $\operatorname{ACIV}$s of the video-level features did not change much compared to the $\operatorname{ACIV}$s over human-only predictions, indicating that simply displaying AI’s prediction did not help humans exploit the available information. Lastly, we demonstrate how ILIV-SHAP explanations surface unique insight into information value that humans lack relative to conventional SHAP for a recidivism prediction task (Angwin et al., 2022; Dressel and Farid, 2018). Detailed descriptions and results of all demonstrations are in Appendices E, D and F.
5 Experiment
We use a preregistered between-subjects online experiment to answer two questions:
  1. 1.
    Can ACIV identify which AI model will result in the best human-AI decision-making?
  2. 2.
    Does ILIV-SHAP improved human-AI decision-making over SHAP alone?
Our experiment assigns participants to one of two AI models with varying ex-ante ACIV and one of three explanation conditions: ILIV-SHAP and SHAP, SHAP, and no explanation. We hypothesize that the higher ACIV model will lead to better human-AI decisions and presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
Metric AI1 AI2
MAPE 14.30% 14.51%
R2 0.81 0.81
$\operatorname{ACIV}$
(in MSE)
$6.5\times 10^{8}$ $3.7\times 10^{8}$
$\operatorname{ACIV}$
(in MAPE)
4.61% 2.00%
Figure 1: Left: AI performance and ACIV. Right: Posterior predicted reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals.
Task and data.
We study an AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al. (2022); Poursabzi-Sangdeh et al. (2021); Holstein et al. (2023)) . This task does not require any domain knowledge, making it suitable for a broad participant population. Additionally, the task is complex enough to benefit from AI assistance. We use the Ames, Iowa Housing Dataset (De Cock, 2011).
Participants and procedure.
We recruited 421 US-based participants via the crowdsourcing platform Prolific. Each participant was randomly assigned to 12 houses in the Ames Iowa Housing Dataset and performed a sequence of 24 trials (where two decisions were elicited for each house, one without AI assistance and one with), split evenly across a human-alone prediction phase and a human-AI prediction phase. In the first 12 trials, participants were asked to predict the house price of the assigned house using six features of the house, including year built, the living area above ground in square feet, the size of garage in car capacity, the number of fireplaces, year remodeled, and ratings of the house’s material and finish. We intentionally reduced the interpretability of the year remodeled and material and finish by witholding what variables they represented and rescaling their values. In the second 12 trials, participants were asked to revise their guesses in the first round after seeing the AI’s prediction and explanation. We incentivized participants with a base reward of $3.00$ USD and a bonus based on the mean squared error (MSE) between their guesses and the true house price in the 24 trials: $\textit{bonus}=3.00-\frac{\textit{MSE}}{3\times 10^{9}}$.
AI models and explanations.
Participants were assigned to one of six experimental conditions resulting from crossing the AI models (2 levels, where AI1 has high $\operatorname{ACIV}$ and AI2 has low $\operatorname{ACIV}$) with the explanation type (3 levels: ILIV-SHAP + SHAP, SHAP, and no explanation). We designed AI1 and AI2 so as to vary their potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features. On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add unique information over human judgments. We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1), so that model accuracy differences would not confound our results. In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction. To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same set-up prior to running the official study. We present ILIV-SHAP explanations of the six features that ranked and highlighted the features where the human-complementary information the AI’s prediction gets from the feature is higher than a threshold (see Appendix I for explanation examples).
Evaluation metrics.
We report the $\operatorname{ACIV}$ of the AI models in the mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not. and the mean absolute percentage error (MAPE). To measure the human unassisted and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model [^4] https://aspredicted.org/jfm6-rrs3.pdf to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$ using R’s brms package (Bürkner, 2017).
[…]
$\nu$ $\sim\text{Gamma}(2,0.1)$
[…]
Results.
Figure 1 confirms the expected ranking of AI models by $\operatorname{ACIV}$. AI1, which had access to the two human-uninterpretable features, achieves higher $\operatorname{ACIV}$ than AI2, which only had access to the four human-interpretable features. Note the similar predictive performance of the two AI models.
Posterior predictions of human-AI team performance according to our regression model also align with our expectations based on the complementary information each model provides (Figure 1 far right). Collapsing across explanation conditions, AI1 reduces APE by more than AI2 ($6.27\%[5.91\%,6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively). Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively).
Figure 1 also illustrates the effect of the different explanation conditions on the human-AI team performance. When the AI has sufficient complementing information (i.e., AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation, all for model AI1).
@@ 6 Discussion and Limitations
Our decision-theoretic framework quantifies the information value of signals available in a human-AI decision setting over the information value of agent decisions. The basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021), even if human agents deviate considerably from Bayesian rational decision-making. We experimentally demonstrate the power of the framework for analyzing and improving human-AI decision-making. Our proposed ILIV-SHAP explanation improves performance over an existing state-of-the-art explanation strategy. Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al. (2023, 2024)). While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments.
One direction for follow-up work is to develop tests that complement the framework by helping evaluators confirm the information that agents are using. Note that low ACIV of a signal over human decisions is not necessarily sufficient for concluding that particular signals were used by the human. It is often unknown whether a human has other private signals offering equivalent information to a feature being analyzed.
@@ Appendix A The Combinatorial Nature of the Value of Signals
When decision-makers are provided with multiple signals, the signals have the combinatorial property by nature. Acknowledgd by recent works in decision theory and game theory [Chen and Waggoner, 2016], one signal may have no information value by itself, but it can be complementary to other signals to provide information value. For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ are uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$. In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff. Though we did not explicitly observe the complementation between signals in our survey of human-AI decision-making tasks (see the results of deepfake detection in supplementary materials), we want to note that our framework can be extended to consider this complementation between signals. We use the Shapley value [Shapley, 1953] to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal. $\phi$ is the average of the marginal contribution of a signal in every combination with other signals.
$\phi(V)=\frac{1}{n}\sum_{V^{\prime}\subseteq\{\Sigma_{1},\ldots,\Sigma_{n}\}/V}{(n-\lvert V\rvert)\choose\lvert V^{\prime}\rvert}^{-1}(\operatorname{ACIV}^{\pi,S}(V^{\prime}\cup V;D^{b})-\operatorname{ACIV}^{\pi,S}(V^{\prime};D^{b}))$ (10)
@@
Appendix B Estimating the Data-Generating Process from Observations
@@
Suppose the data-generating process is estimated from the empirical distribution of the observed realizations of the payoff state value and the signal, $D=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
$\hat{\pi}_{D}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
In applying our framework, the rational DM’s decision rule $d^{r}$ should not overfit to the observed dataset. This threatens the usefulness of analysis using the DGP. To avoid this situation, we can either collect more observations or “coarsen” the signal space, where “coarsen” means reducing the signal space by aggregating similar signals into a single signal. We propose a systematic procedure for the latter since it may be too costly or otherwise not possible to collect more observations. Our objective is to find the coarsened signal structure that has the largest signal space and avoids the rational decision rule from overlapping. Denote the maximum tolerance factor for overfitting as $\epsilon$. The rational DM’s decision rule $d^{r}_{\hat{\pi}_{D^{\prime}}}$ that is estimated from a set of observations $D^{\prime}$ should perform similarly on $D^{\prime}$ and on the evaluation of a holdout set $D\setminus D^{\prime}$, i.e.,
$\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]-\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D\setminus D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]\leq\epsilon*\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]$ (12)
Algorithm 2 Estimating the data-generating process from a set of observations
1:  Input: Observed dataset $C=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Overfitting tolerance $\epsilon$, Ratio of the holdout set $\delta$.